OpenStudy (anonymous):

Which will result in a perfect square trinomial? (3x – 5)(3x – 5) (3x – 5)(5 – 3x) (3x – 5)(3x + 5) (3x – 5)(–3x – 5)

OpenStudy (anonymous):

@AlexandervonHumboldt2 @FLAMER123

OpenStudy (zale101):

a perfect trinomials is when you have two numbers that are the same and they can add up and at the same time get multiplied. \(\large a^2+2ab+b^2=\large (a+b)^2\)

OpenStudy (zale101):

for example: \(x^2+4x+4\) 2+2=4x 2*2=4

OpenStudy (anonymous):

@Zale101 so What Is The Answer